# Yang-Lee Zeros in Quantum Phase Transition: An Entanglement Perspective

### v1

@@ Abstract
Here we define $\cosh\phi:=\Delta$ and $\exp(ik_{j})=\sin\frac{\phi}{2}(x_{j}+i)/\sin\frac{\phi}{2}(x_{j}-i)$, where $k_{j}$ is the quasimomentum. The index is given by $j=1,\cdots,M$ with $M$ being the number of the excited magnons. If we take the limit $\delta:=\Delta-1\rightarrow 0$, we can approximate $\phi$ as $\phi\simeq\sqrt{2\delta}$. When $\phi=0$ ($\Delta$ = 1), Eq. (S.65) can be reduced as
$\left(\frac{x_{j}+i}{x_{j}-i}\right)^{L}=\prod_{l\neq j}\left(\frac{x_{j}-x_{l}+2i}{x_{j}-x_{l}-2i}\right),$ (S.66)
@@
which consists of the Bethe ansatz equations of the Heisenberg model. In this case the variable $x_{j}$ can be rewritten as $x_{j}=\cot(k_{j}/2)$. From Eq. (S.66), we can see $x_{j}=\infty(k_{j}=0)$ is a set of solutions of the Bethe ansatz, which represent all magnons have zero momentum or equal-weight superposition on each site of the system. These states are ground states of the Heisenberg model deduced from the energy:
$E_{M}=E_{0}+\sum_{j=1}^{M}J(1-\cos k_{j})=E_{0}+\sum_{j=1}^{M}\frac{2J}{x_{j}^{2}+1}.$ (S.67)
@@
For the Heisenberg model with a system size $L$, there is $(L+1)$-degeneracy of the ground state, originating from the SU(2) symmetry of the model. These eigenstates can be expressed as $|M\rangle=(S^{-})^{M}|\Omega\rangle$ where $S^{-}=\sum_{j}S_{j}^{-}$. However, when we have a finite $\phi$ for anisotropy, the degeneracy will be lifted. To be specific, we begin from the expansion of the Bethe ansatz equations (S.65) of the XXZ model. Since we only consider the distribution of Yang-Lee zeros around the phase transition point $\Delta=1$, we take the limit $\phi\rightarrow 0$ and expand the equation as
$\left(\frac{\sin\frac{\phi}{2}x_{j}+i\frac{\phi}{2}\cos\frac{\phi}{2}x_{j}}{\sin\frac{\phi}{2}x_{j}-i\frac{\phi}{2}\cos\frac{\phi}{2}x_{j}}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l})+i\phi\cos\frac{\phi}{2}(x_{j}-x_{l})}{\sin\frac{\phi}{2}(x_{j}-x_{l})-i\phi\cos\frac{\phi}{2}(x_{j}-x_{l})}.$ (S.68)
@@
$L\cot\frac{\phi}{2}x_{j}=2\sum_{l\neq j}\cot\frac{\phi}{2}(x_{j}-x_{l}),$ (S.69)
which is referred to as simplified Bethe ansatz equations (SBAE). If we define $y_{j}=\frac{\phi}{2}x_{j}$, the SBAE can be also rewritten as
$L\cot y_{j}=2\sum_{l\neq j}\cot(y_{j}-y_{l}).$ (S.70)
We can see this equation is purely an algebraic equation on $\cot y_{j}$ where the solutions only depend on the size $L$ and the magnon number $M$. Meanwhile, the energy of the $M$-magnon excitation is given by
$E=E_{0}+\sum_{j=1}^{M}\frac{J\sinh^{2}\phi}{\cosh\phi-\cos\phi x_{j}}$ $=$ $E_{0}+\sum_{j=1}^{M}\frac{2J\delta}{1+\delta-\cos\phi x_{j}}$ (S.71)
@@
In the expansion above, we first assume that $\cot y_{j}$ is finite so that we can expand the Bethe ansatz equation (S.65) up to the first order of $\phi$. Now we can also see this point from the expression (S.71). Divergence of the factor $\cot y_{j}$ will increase $2J$ to the energy, indicating that this solution does not originate from the original $(L+1)$-degeneracy of the Heisenberg spin chain. Therefore, we can exactly obtain all the deviated ground states from the SBAE. The finiteness of $\cot y_{j}$ of deviated ground states can be also seen from the SBAE (S.69). We assume that there are divergences of $\cot y_{j}$ first. Then there must be a series of $\cot y_{j}$ which satisfy $\cot y_{j}=a_{j}N$ where $j=1,\cdots,m$ with $N$ being a divergent number and $a_{j}$ being a constant. Since some of the solutions $\cot y_{j}$ may not diverge, we always have $m\leqslant M\leqslant L$. Then the SBAE becomes
$La_{j}=2\sum_{l\neq j}\frac{a_{j}a_{l}}{a_{l}-a_{j}}.$ (S.72)
@@
$L\zeta_{j}=2\sum_{l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}.$ (S.74)
@@
$\sum_{j}\zeta_{j}^{2}$ $=$ $\frac{2}{L}\sum_{j,l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}\zeta_{j}$ (S.75)
@@
where we exchange the label $l,j$ in the second equality. Hence, by summation of the right-hand side of the two equality, we obtain
$\sum_{j}\zeta_{j}^{2}$ $=$ $-\frac{1}{L}\sum_{j,l\neq j}(1+\zeta_{l}\zeta_{j})$ (S.76)
@@
where we use $\sum_{j}\zeta_{j}=0$ from Eq. (S.74). Hence, the summation of $\zeta_{j}^{2}$ can be simplified as
@@
By revisiting the expression of energy (S.71), we have
$E=E_{0}+\sum_{j=1}^{M}J\delta(1+\cot^{2}y_{j})=E_{0}+J\delta\left(M+\sum_{j=1}^{M}\zeta_{j}^{2}\right)=E_{0}+J\delta\frac{M(L-M)}{L-1},$ (S.77)
which can indicate that the energy reaches the maximum for $M=\frac{L}{2}\left(\frac{L-1}{2}\right)$ when $L$ is even (odd). With the expression of energy (S.77), we show the expression of the partition function at absolute zero as
$Z=\lim_{T\rightarrow 0}\text{Tr}e^{-\beta H}=\lim_{T\rightarrow 0}\sum_{M=0}^{L}e^{-\beta E_{M}}=\lim_{T\rightarrow 0}e^{-\beta E_{0}}\sum_{M=0}^{L}z^{M(L-M)}.$ (S.78)
@@
Finally, we discuss the Yang-Lee edge singularity associated with the Yang-Lee zeros. Around the phase transition point $\Delta=1$, we consider the susceptibility Takahashi (1999)
$\chi=\frac{2s_{z}}{h}=\frac{4\gamma}{J\pi(\pi-\gamma)\sin\gamma}[1+O(h^{2})+O(h^{4\gamma/(\pi-\gamma)})],$ (S.80)
@@
$\chi(h=0)\simeq\frac{4\gamma}{J\pi\left(-i\sqrt{2\delta}\right)i\sqrt{2\delta}}=\frac{2}{J\delta}\propto\delta^{-1}.$ (S.82)
@@
$\Delta E=\text{Re}[J\delta],$ (S.83)
which is determined by the Bethe ansatz. Hence, from the solution, we can see the critical exponent $\varphi=1$, where we define $\Delta E=J\times\text{Re}[\delta]^{\varphi}$.

### v2

@@ Appendix B Entanglement Entropy in Non-interacting Fermionic Systems
Here we define $\cosh\phi:=\Delta$ and $\exp(ik_{j})=\sin\frac{\phi}{2}(x_{j}+i)/\sin\frac{\phi}{2}(x_{j}-i)$, where $k_{j}$ is the quasimomentum. The index is given by $j=1,\cdots,M$ with $M$ being the number of the excited magnons. If we take the limit $\delta:=\Delta-1\rightarrow 0$, we can approximate $\phi$ as $\phi\simeq\sqrt{2\delta}$. When $\phi=0$ ($\Delta$ = 1), Eq. (78) can be reduced as
$\left(\frac{x_{j}+i}{x_{j}-i}\right)^{L}=\prod_{l\neq j}\left(\frac{x_{j}-x_{l}+2i}{x_{j}-x_{l}-2i}\right),$ (79)
@@
which consists of the Bethe ansatz equations of the Heisenberg model. In this case the variable $x_{j}$ can be rewritten as $x_{j}=\cot(k_{j}/2)$. From Eq. (79), we can see that $x_{j}=\infty(k_{j}=0)$ is a set of solutions of the Bethe ansatz, which represents that all magnons have zero momentum or equal-weight superposition on each site of the system. These states are ground states of the Heisenberg model deduced from the energy:
$E_{M}=E_{0}+\sum_{j=1}^{M}J(1-\cos k_{j})=E_{0}+\sum_{j=1}^{M}\frac{2J}{x_{j}^{2}+1}.$ (80)
@@
For the Heisenberg model with a system size $L$, there is $(L+1)$-degeneracy of the ground state, originating from the SU(2) symmetry of the model. These eigenstates can be expressed as $|M\rangle=(S^{-})^{M}|\Omega\rangle$ where $S^{-}=\sum_{j}S_{j}^{-}$. However, when we have a finite $\phi$ for anisotropy, the degeneracy will be lifted. To be specific, we begin from the expansion of the Bethe ansatz equations (78) of the XXZ model. Since we only consider the distribution of Yang-Lee zeros around the phase transition point $\Delta=1$, we take the limit $\phi\rightarrow 0$ and expand the equation as
$\left(\frac{\sin\frac{\phi}{2}x_{j}+i\frac{\phi}{2}\cos\frac{\phi}{2}x_{j}}{\sin\frac{\phi}{2}x_{j}-i\frac{\phi}{2}\cos\frac{\phi}{2}x_{j}}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l})+i\phi\cos\frac{\phi}{2}(x_{j}-x_{l})}{\sin\frac{\phi}{2}(x_{j}-x_{l})-i\phi\cos\frac{\phi}{2}(x_{j}-x_{l})}.$ (81)
@@
$L\cot\frac{\phi}{2}x_{j}=2\sum_{l\neq j}\cot\frac{\phi}{2}(x_{j}-x_{l}),$ (82)
which is referred to as simplified Bethe ansatz equations (SBAEs). If we define $y_{j}=\frac{\phi}{2}x_{j}$, the SBAEs can be rewritten as
$L\cot y_{j}=2\sum_{l\neq j}\cot(y_{j}-y_{l}).$ (83)
We can see that this equation is purely an algebraic equation on $\cot y_{j}$, where the solutions only depend on the size $L$ and the magnon number $M$. Meanwhile, the energy of the $M$-magnon excitation is given by
$E=E_{0}+\sum_{j=1}^{M}\frac{J\sinh^{2}\phi}{\cosh\phi-\cos\phi x_{j}}$ $=$ $E_{0}+\sum_{j=1}^{M}\frac{2J\delta}{1+\delta-\cos\phi x_{j}}$ (84)
@@
In the expansion above, we first assume that $\cot y_{j}$ is finite so that we can expand the Bethe ansatz equation (78) up to the first order of $\phi$. Now we can also see this point from the expression (84). Divergence of the factor $\cot y_{j}$ will increase $2J$ to the energy, indicating that this solution does not originate from the original $(L+1)$-degeneracy of the Heisenberg spin chain. Therefore, we can exactly obtain all the deviated ground states from the SBAE. The finiteness of $\cot y_{j}$ of the deviated ground states can also be seen from the SBAE (82). We assume that there are divergences of $\cot y_{j}$ first. Then there must be a series of $\cot y_{j}$ which satisfy $\cot y_{j}=a_{j}N$, where $j=1,\cdots,m$, with $N$ being a divergent number and $a_{j}$ being a constant. Since some of the solutions $\cot y_{j}$ may not diverge, we always have $m\leqslant M\leqslant L$. Then the SBAE becomes
$La_{j}=2\sum_{l\neq j}\frac{a_{j}a_{l}}{a_{l}-a_{j}}.$ (85)
@@
$L\zeta_{j}=2\sum_{l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}.$ (87)
@@
$\sum_{j}\zeta_{j}^{2}$ $=$ $\frac{2}{L}\sum_{j,l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}\zeta_{j}$ (88)
@@
where we exchange the label $l,j$ in the second equality. Hence, by summation of the right-hand side of the two equalities, we obtain
$\sum_{j}\zeta_{j}^{2}$ $=$ $-\frac{1}{L}\sum_{j,l\neq j}(1+\zeta_{l}\zeta_{j})$ (89)
@@
where we use $\sum_{j}\zeta_{j}=0$ from Eq. (87). Hence, the summation of $\zeta_{j}^{2}$ can be simplified as
@@
By revisiting the expression of energy (84), we have
$E=E_{0}+\sum_{j=1}^{M}J\delta(1+\cot^{2}y_{j})=E_{0}+J\delta\left(M+\sum_{j=1}^{M}\zeta_{j}^{2}\right)=E_{0}+J\delta\frac{M(L-M)}{L-1},$ (90)
which can indicate that the energy reaches the maximum for $M=\frac{L}{2}\left(\frac{L-1}{2}\right)$ when $L$ is even (odd). With the expression of energy (90), we show the expression of the partition function at absolute zero as
$Z=\lim_{T\rightarrow 0}\text{Tr}e^{-\beta H}=\lim_{T\rightarrow 0}\sum_{M=0}^{L}e^{-\beta E_{M}}=\lim_{T\rightarrow 0}e^{-\beta E_{0}}\sum_{M=0}^{L}z^{M(L-M)}.$ (91)
@@
Finally, we discuss the Yang-Lee edge singularity associated with the Yang-Lee zeros. Around the phase transition point $\Delta=1$, we consider the susceptibility [60]
$\chi=\frac{2s_{z}}{h}=\frac{4\gamma}{J\pi(\pi-\gamma)\sin\gamma}[1+O(h^{2})+O(h^{4\gamma/(\pi-\gamma)})],$ (93)
@@
$\chi(h=0)\simeq\frac{4\gamma}{J\pi\left(-i\sqrt{2\delta}\right)i\sqrt{2\delta}}=\frac{2}{J\delta}\propto\delta^{-1}.$ (95)
@@
$\Delta E=\text{Re}[J\delta],$ (96)
which is determined by the Bethe ansatz. Hence, from the solution we can see the critical exponent $\varphi=1$, where we define $\Delta E=J\times\text{Re}[\delta]^{\varphi}$.